Good day,
Ammonia, NH3, is a weak base with a Kb value of 1.8×10−5
I calculated The pH of a 0.425 M ammonium solution is 11.44
What is the percent ionization of ammonia at this concentration?
Express your answer with the appropriate units.
Thanks
1
Expert's answer
2020-08-11T12:47:32-0400

The percent ionization of the ammonium ion is calculated by estimating the fraction of ammonia dissociated from the initial concentration. The percentage ionization has been calculated as follows:

Percent ionization = ([NH3]dissociated / [NH3]initial)×100

[NH3]initial = 0.425 mol/L

[NH3]dissociated = [NH4+] = [OH-]

pH+pOH=14

pH=11.44

11.44 + pOH = 14

pOH= 2.56

pOH=−log[OH−]

2.56 = −log[OH−]

10−(2.56) = [OH−]

[OH−] = 10−(2.56) = 0.00275 mol/L

Percent ionization = (0.00275 / 0.425)×100 = 0.64 %

Answer: 0.64 %